right wrong
Pick an answer to see how members did
Log this miss in the free GMAT error log
Get the error logGet one like this every day: the free 60-day plan
Start the plan(A) 0.1
(B) 0.2
(C) 0.25
(D) 0.4
(E) 0.6
OA A
Source: Manhattan Prep
Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.
See classesLearn live with an expert or move at your own pace. Every option includes the complete TTP study system.
Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.
with Josh Braslow, 20+ years in test prep
Self-paced EA prep. Study on your schedule.
with
Complete access from day one. Study on your schedule.
right wrong
Pick an answer to see how members did
Log this miss in the free GMAT error log
Get the error logGet one like this every day: the free 60-day plan
Start the planApproach 1;BTGmoderatorDC wrote:From a group of 5 managers (Joon, Kendra, Lee, Marnie and Noomi), 2 people are randomly selected to attend a conference in Las Vegas. What is the probability that Marnie and Noomi are both selected?
(A) 0.1
(B) 0.2
(C) 0.25
(D) 0.4
(E) 0.6
Our goal is to find P(M and N both selected)BTGmoderatorDC wrote:From a group of 5 managers (Joon, Kendra, Lee, Marnie and Noomi), 2 people are randomly selected to attend a conference in Las Vegas. What is the probability that Marnie and Noomi are both selected?
(A) 0.1
(B) 0.2
(C) 0.25
(D) 0.4
(E) 0.6
OA A
Source: Manhattan Prep
There are a total of 5C2 = 5!/(3! x 2!) = (5 x 4)/2 = 10 ways of choosing two people from a group of five people. Since the selection of Marnie and Noomi corresponds to one of these choices, the probability of this selection is 1/10 = 0.1.BTGmoderatorDC wrote:From a group of 5 managers (Joon, Kendra, Lee, Marnie and Noomi), 2 people are randomly selected to attend a conference in Las Vegas. What is the probability that Marnie and Noomi are both selected?
(A) 0.1
(B) 0.2
(C) 0.25
(D) 0.4
(E) 0.6
OA A
Source: Manhattan Prep
